A propeller driven plane leaves the airport and travels north at 200 mph. Two hours later, a jet plane leaves the same airport and travels north at 600 mph. How long will it take the jet plane to catch up to the propeller driven plane?
Answer by stanbon(75887)

A propeller driven plane leaves the airport and travels north at 200 mph. Two hours later, a jet plane leaves the same airport and travels north at 600 mph. How long will it take the jet plane to catch up to the propeller driven plane?
--------------
Prop plane DATA:
Rate = 200 mph ; distance = x miles ; time = d/r = x/200 miles
---------
Jet plane DATA:
Rate = 600 mph ; distance = x miles ; time = d/r = x/600 miles
------------
EQUATION:
jet time + 2 hr. = plane time
x/600 + 2 = x/200
Multiply thru by 600 to get:
x + 1200 = 3x
2x=1200= 600 miles
Time for ject plane: 600/x = 600/600 = 1 hour.
Cheers,
Stan H.